OpenStudy (anonymous):

The Big Screamer Coaster carries 92 people altogether. Some of its cars carry 4 passengers and the rest carry 6 passengers. There are 3 less 6-passenger cars the there are 4-passenger cars. Write and equation to solve how many 4-passenger cars there are. @aadsadsf

OpenStudy (anonymous):

1. Let's call the number of 4-passenger cars "P"... 2. You know that there are 3 less 6-passenger cars than 4-passenger cars. So there are (P - 3) 6-passenger cars. 3. The number of people from all 4-passenger cars is just 4 x P.... 4 people per car times the number of 4-passenger cars. 4. In the same way as in step 3, the number of people from all 6-passenger cars is: 6 x (P-3) remembering that the number of 6-passenger cars is P-3 (from step 2). 5. The total number of people from both cars will be step + step 4... 4P + 6(P-3), and the problem says that this number is 92. So 4P + 6(P-3) = 92 6. Simplify the equation in step 5. 7. Solve the simplified equation for P, the number of 4-passenger cars.

OpenStudy (anonymous):

In your problem, you get lucky (I guess...) because they don't even require you to solve for P... they just want to see the simplified equation you find in step 6.

OpenStudy (anonymous):

omgg thankk you soooo muchh!

OpenStudy (anonymous):

Yeah what he said lol

OpenStudy (anonymous):

Please read through it and make sure you understand how I set it up... it's not too tough once you learn how to "translate" English sentences into "math sentences" (meaning equations).

OpenStudy (anonymous):

Let me know if there's any part or step you don't get :)

OpenStudy (anonymous):

ok i understand it:) thank you for all your helpp

OpenStudy (anonymous):

And in future problems, use a similar process... set a variable for whatever you want to solve, then write the expressions to match the problem's sentences, then simplify, and finally solve. Glad to help :)

OpenStudy (anonymous):

im not sure what p would be woulf it be 7 and 2/5

OpenStudy (anonymous):

Remember, the question is only asking for the equation, not the actual number for P. But in your solving for P, I think you might have made an error somewhere. Here's what I get... From step 5.... 4P + 6(P-3) = 92 Simplifying... 4P + 6P - 18 = 92 10P - 18 = 92 10P = 92 + 18 10P = 110 10P/10 = 110/10 P = 11

OpenStudy (anonymous):

Oh sorry!! the question DOES want you to solve for P... oops!! Set up the equation AND solve... sorry I missed that before.

OpenStudy (anonymous):

ohh ok thankss